# Large Language Models can be Guided to Evade AI-Generated Text Detection

### v5

@@ 5.3 Examples of the generated text
Object: Sushi Station Keywords: revolving sushi bar, affordable, krab mix, creamy sauce, fresh menu. Sentiment: Positive
$\mathcal{P}_{\text{AI}}$
@@
AI: If you’re a sushi lover, you have to check out Sushi Station! This revolving sushi bar is not only affordable, but it offers a fresh and delicious menu that will leave you craving more. My personal favorite is the krab mix sushi with its creamy sauce - it’s simply heavenly. The staff is friendly and attentive, and the atmosphere is perfect for a casual night out with friends or family. Don’t hesitate to give Sushi Station a try - you won’t be disappointed!
0.99
@@
SICO-Para: Sushi enthusiasts, please take note: Sushi Station provides an unbeatable revolving sushi bar experience at wallet-friendly prices. Their menu is of the freshest quality and tastes nothing short of delicious. My top pick has to be the krab mix sushi; its creamy sauce is pure bliss. Additionally, their friendly staff provided excellent service, while the casual ambiance proved ideal for a relaxed night out with friends or family. Don’t hesitate to give Sushi Station a try, you won’t regret it!
0.00
@@
SICO-Gen: I recently tried out Sushi Station’s revolving sushi bar and I must say, it was an amazing experience. The affordable pricing along with the wonderful krab mix and creamy sauce they offer is what really made it stand out from the rest. The menu was constantly updated with fresh and delicious options making it hard to choose just one. If you’re looking for a great sushi experience, Sushi Station is definitely worth a visit.
0.04
@@ 6 Conclusion
In conclusion, we have proposed a novel in-context learning approach, SICO, designed to guide LLMs in generating text that can effectively evade detectors. Our extensive experiments on evasion demonstrate the superior performance of SICO, which significantly reduces the detection capabilities of existing AI text detectors across three tasks. A comprehensive human evaluation shows SICO text can achieve human-level readability and task completion rates. The experiment in the wild shows the functionality of SICO in real life.
@@ 7 Ethics statement
The intention of this paper is not to offer a potential method for evading AI-generated text detection systems. Instead, our aim is to raise awareness within the broader community about the vulnerabilities of existing AI-generated text detection systems to such technology. As many LLMs are public available and free to use, many people can adjust their prompt and generate text that evades these detectors. Given the ease of evasion illustrated in this study, these detectors are not robust yet.
We hope the research community can stress test their detectors against text generated by carefully crafted prompt, and create more robust detectors in the future. To support the research in this field, we will make our training methods and relevant data/code publicly available.
@@ A.1.3 Word Substitution
We employ WordNet synsets to derive synonyms for given words. During optimization of in-context examples, we only substitute content words, namely nouns, verbs, adjectives, and adverbs. Furthermore, we part-of-speech (POS) tag of the synonym to ensure it aligns with the original word. For POS tagging, we utilize the Stanford POS Tagger Toutanova et al. [2003]. Additionally, to maintain fluency in the modified text after substitution, we employed a pretrained mask language model to exclude synonyms with low likelihood. In experiment we use RoBERTa-base model [Liu et al., 2019].
@@
Task
Task instruction $p_{\text{task}}$
Writing
Based on the description, complete an academic paragraph in human style writings:
@@
QA
Based on the description, answer questions in human style writings:
@@
Review
Based on the description, write a human review about the given object and keywords, with a specified sentiment:
@@
Paraphrase
Based on the description, rewrite this in human-style writings:
@@ A.2.1 DIPPER
We choose the best evasion performance parameter setting from the original paper [Krishna et al., 2023], which is 60 for lexical diversity and 60 for re-ordering. And we set sampling temperature to 0.75, following the original implementation.
@@ A.2.2 Human prompt
We carefully design a paraphrase prompt based on the detection idea of GPTzero [Tian, 2023], which distinguish the AI-generated content from Human-written by perplexity and burstiness, stated by its creator [^3] https://theconversation.com/we-pitted-chatgpt-against-tools-for-detecting-ai-written-text-and-the-results-are-troubling-199774. Perplexity is the concept raised in NLP field, which measures how well a language model predicts a text sample. A text with a lower perplexity score indicates that the language model is better at calculating the next word that is likely to occur in a given sequence. On the other hand, burtiness basically measures the variation between sentences, including sentence length and structures. The lower the values for these two factors, the more likely it is that a text was produced by an AI. Table 10 shows the prompt we designed.
@@ Appendix B Extra Experiments
B.1 Detection capability of human
We conducted a experiment of real-life human detection rate. We randomly selected 150 examples (50 each from AI, SICO-Para, and human responses) from QA task. Four human annotators were then asked to identify whether each answer was AI-generated or human-written. Table 11 shows that SICO remarkably reduces the probability of being recognized by humans. Additionally, the data also reveals that human detection capabilities are not highly accurate compared with machine detectors. Approximately half of AI-generated texts went undetected, and 10% of human-written texts were mistakenly identified as AI-generated.
Table 11: Detection rate from human anotators. ’AI %’ indicates the proportion of texts identified as AI-generated by annotators.
Text Source AI %
AI 54%
SICO-Para 26%
Human 10%
B.2 Semantic preserving
We measure semantic similarity using t5-based sentence encoder [Ni et al., 2022], which leads in semantic text similarity task of MTEB benchmark [Muennighoff et al., 2023]. Table 12 reports a comparison of the cosine similarity of text before and after paraphrasing by different methods.
@@ B.2 Semantic preserving
B.3 Consistent effectiveness of SICO
SICO consistently procures optimal detection evasion prompts, irrespective of the diversity in AI-human text pairs and randomized samples drawn from the LLMs. This demonstrates the effectiveness of SICO in various initial conditions and settings, highlighting its applicability to diverse scenarios. Figure 4 presents the AI probability of the paraphrased training data, generated by LLM using the current best prompt, denoted as $\mathcal{U}(p^{*})$ in Equation 1. SICO successfully optimizes the initial prompt (at step 0) and converges to a low level in three separate trials across all three tasks, with different AI-human text pairs.
@@ B.3 Consistent effectiveness of SICO
B.4 SICO performs better against more capable detectors
@@ Appendix C Detectors
GPT-3.5 detector is trained on Human ChatGPT Comparison Corpus (HC3) dataset [Guo et al., 2023], which including answers generated by ChatGPT and human. English-version of HC3 dataset contains five splits: reddit-eli5, OpenQA, Wikipedia, medicine and finance. The base model is RoBERTa-base and we use the model that only take answers as input.
@@ C.3 DetectGPT
We use z-score implementation of DetectGPT and set sample number to 100 and replacement ratio to 0.3. The inner language model is GPT2-medium and the mask language model is t5-large [Raffel et al., 2020].
@@ C.6 OpenAI Detector
OpenAI detector is officially provided by OpenAI after the release of ChatGPT. Although it only offers a web interface, we adopted the API implementation from [Krishna et al., 2023], which uses “model-detect-v2” in the OpenAI API. Through reverse engineering of the website, we determined that the web interfaces indeed call this model.
@@ Appendix D Human Evaluation Details
For each text, we show two questions for the human annotator. In terms of readability, we present the human annotator five options, with the scale of 1 to 5. The higher the value, the more readable of the presented texts. The question is identical for three tasks. The actual question and options are presented in Table 13. For task completion rate, we design three task-specific questions, as show in Table 14. Figure 6 shows the interface of our annotation platform. We estimated that the evaluation time of each text ranges from 60 to 90 seconds.
@@
Appendix E Details of Real-life Experiments
We employ the Praw package, a Reddit API wrapper, to write the auto-reply bot. On each run, the bot first collect the posts that are newly submitted within 2 hours and extract the title of poster as question. Then the bot will utilize GPT-3.5 to generate the original answer, and use SICO-Para to paraphrase it. And we set rules to filter out personal question like “Which movie do you like best?”. Because GPT-3.5 will reply like “I’m sorry, as a language model…”, which will directly explore itself. The prompt we deploy is trained for QA task. We only reply in the communities that do not ban AI-generated content: specifically explainlikeimfive and NoStupidQuestions. We only generate 40 responses and deleted them after collecting the data.
Appendix F Datasets
@@
Rev-Gen
Write a review about given object and key words, with specified sentiment:
@@
Appendix G Evade Watermarking Detection
SICO-Para can also be utilized to evade watermark detection, similarly to paraphrase approach. The watermarking algorithm we applied was introduced by Kirchenbauer et al. [2023], which only requires access to the LLM’s logits at each time step to add watermarks. This algorithm operates in three steps:
@@ Appendix G Evade Watermarking Detection
In our experiments, we used the text generated by a watermarked GPT-2, provided by Krishna et al. [2023]. We employed the GPT-3.5 detector as proxy detector for training. The AUC and and the detection accuracy associated with various paraphrasing methods are presented in Table 18, where the threshold is set to 2.2 for accuracy measurement.
The results reveal that SICO-Para significantly outperforms other paraphrase techniques in evading watermark detection. Notably, both the AUC score and detection accuracy of SICO-Para are significantly lower than that of other methods. Furthermore, even the simpler baseline model, GPT-Para, proved to be effective at evading the watermark detection.
@@
Appendix H Evasion Performance
H.1 ROC Curves
@@ H.2 Detection Accuracy
Task
Feature
Writing
Human tend to provide more specific details and facts, often including dates, numbers, and names. They also tend to use more complex sentence structures, including semicolons and parenthetical phrases, to convey additional information. However, their writing remains clear and concise, and they avoid unnecessary repetition or wordiness.
QA
Human’s writing style is characterized by the provision of specific and technical information. They delve into the science behind a topic or the details of a process. They also challenges or adds nuance to the information, providing alternative explanations or clarifying misconceptions.
Review
Human writings are generally shorter and more to-the-point. They often use bullet points or numbered lists to convey their thoughts. They tend to be more critical and focused on their own personal experiences. They also use informal language and occasionally include personal anecdotes or opinions.
Appendix I Examples of Text Generated by SICO
The examples of text generated by SICO across three tasks are presented in Tabel 21-23.
@@ Appendix I Examples of Text Generated by SICO
Question: Why does changing screen resolution (more specifically, a custom resolution) will harm your computer?
$\mathcal{P}_{\text{AI}}$
AI: Changing the screen resolution, including a custom resolution, will not harm your computer. It may result in display issues or reduced image quality if the resolution is not compatible with your monitor’s capabilities. However, this can be resolved by simply reverting to the previous resolution or finding a suitable one.
0.99
SICO-Para: Don’t worry, changing your screen resolution won’t harm your computer, even if you decide to go custom. The only trouble is that it may not work well with your monitor and cause image problems or lesser quality. Easy solution though - just go back to the old resolution or find a matching one.
0.00
SICO-Gen: Using custom screen resolutions places a greater burden on system processing units, potentially leading to increased strain and overheating. Moreover, setting a resolution to unsupported display dimensions can put undue strain on graphics hardware, possibly causing irreparable damage. Adjusting to custom resolutions might, therefore, jeopardize the computer, leading to hardware malfunctions, degraded display output quality, and issues with other essential system components.
0.00
Question: How do they calculate how much calories you’ve burned after an exercise?
$\mathcal{P}_{\text{AI}}$
AI: The number of calories burned during exercise is calculated by various factors such as body weight, the intensity of the exercise, the duration of the activity, and the person’s metabolic rate. Some popular fitness trackers and machines use algorithms and formulas based on these factors to estimate the number of calories burned. However, these numbers are only estimates and can vary depending on individual factors such as muscle mass, fitness level, and metabolic rate.
0.99
SICO-Para: If you wanna know how many calories you burn during exercise, there are a few things to keep in mind. Your weight, the intensity of your workout, and how long you’re doing it for all come into play, as does your metabolic rate. Popular fitness trackers and machines use fancy calculations based on all these factors to give you an estimate of how many cals you burned. But, listen up, folks - these are only estimates! Your muscle mass, fitness, and metabolism are gonna differ from person to person, so results will vary. Clear?
0.05
SICO-Gen: The calculation of calorie consumption requires assessment of parameters such as exercise type, duration, and intensity, as well as the individual’s body measurements. Several formulas exist to compute the total number of calories burned, including the Harris-Benedict equation, which computes resting metabolic rate and establishes parameters for calorie expenditure. Modern heart rate monitors and wearable fitness trackers utilise the latest algorithms to compute calorie consumption by taking into account the heart rate, age, and weight of the user.
0.00
@@
Object: hotel in Las Vegas Keywords: Penthouse, staff, casino, restaurants, excellent experience. Sentiment: Positive
$\mathcal{P}_{\text{AI}}$
AI: I had an amazing experience staying at a penthouse in Las Vegas! The staff was incredibly friendly and accommodating, always going above and beyond to make my stay comfortable and enjoyable. The casino was also a great addition, providing endless entertainment and excitement. And when it came to dining options, the variety of restaurants available was impressive. Overall, staying at this hotel was an excellent experience that I won’t forget anytime soon. I highly recommend it to anyone looking for a luxurious and unforgettable Las Vegas getaway.
0.75
SICO-Para: Had an amazing stay at a penthouse in Las Vegas. The staff were incredibly friendly and made my stay comfortable and enjoyable. Casino provided endless entertainment and variety of restaurants impressive. Overall, excellent experience and highly recommend for luxurious and unforgettable Las Vegas getaway.
0.03
SICO-Gen: I had the most incredible experience staying at a luxurious penthouse in a Las Vegas hotel. The staff was exceptional and provided top service. The casino and restaurants on the premises were outstanding, making my stay all the more enjoyable. Overall, I highly recommend this hotel for an excellent experience in sin city.
0.06
Object: Tia Rosa restaurant Keywords: Mexican food, spicy, bland taste, sweet syrup, tacos. Sentiment: Negative
$\mathcal{P}_{\text{AI}}$
AI: I recently visited Tia Rosa restaurant hoping to enjoy some authentic Mexican food. Unfortunately, my experience was far from satisfactory. The food was overly spicy, to the point where it overwhelmed any other flavors that may have been present. On the other hand, the tacos were disappointingly bland and lacked any real taste. To add insult to injury, the sweet syrup that was served with my meal just didn’t pair well with the savory dishes. Overall, my visit to Tia Rosa left me feeling disappointed and unsatisfied.
0.84
SICO-Para: Recently went to Tia Rosa restaurant looking for authentic Mexican food. However, experience was unsatisfactory. Food was overly spicy, overpowering other flavours. Tacos lacked taste. Sweet syrup served with meal did not complement savoury dishes. Overall left feeling unsatisfied and disappointed.
0.02
SICO-Gen: My visit to Tia Rosa restaurant was a complete letdown. I was expecting some spicy and flavorful Mexican food, but unfortunately, the dishes were quite bland. Even the tacos, which are usually a safe bet, were a disappointment. The only thing worthy of praise is the sweet syrup, but this alone cannot make up for the overall taste. I won’t be returning to Tia Rosa anytime soon.
0.04
@@
Writing Start: The code itself was patterned so that most control codes were together, and all graphic codes were together, for ease of identification. The first two columns (32 positions) were reserved
$\mathcal{P}_{\text{AI}}$
@@
AI: The code itself was patterned so that most control codes were together, and all graphic codes were together, for ease of identification. The first two columns (32 positions) were reserved for control codes, while the remaining 94 columns (94 positions) were designated for graphic codes. This pattern allowed for efficient processing and recognition of the different types of codes. Additionally, the code was designed to be easily modifiable, with the ability to add or remove codes as needed. These features made the code a practical and versatile tool for programming and data processing tasks.
0.99

### v6

@@ 5.3 Examples of SICO text
Object: Sushi Station Keywords: revolving sushi bar, affordable, krab mix, creamy sauce, fresh menu. Sentiment: Positive $\mathcal{P}_{\text{AI}}$
@@
AI: If you’re a sushi lover, you have to check out Sushi Station! This revolving sushi bar is not only affordable, but it offers a fresh and delicious menu that will leave you craving more. My personal favorite is the krab mix sushi with its creamy sauce - it’s simply heavenly. The staff is friendly and attentive, and the atmosphere is perfect for a casual night out with friends or family. Don’t hesitate to give Sushi Station a try - you won’t be disappointed! 0.99
@@
SICO-Para: Sushi enthusiasts, please take note: Sushi Station provides an unbeatable revolving sushi bar experience at wallet-friendly prices. Their menu is of the freshest quality and tastes nothing short of delicious. My top pick has to be the krab mix sushi; its creamy sauce is pure bliss. Additionally, their friendly staff provided excellent service, while the casual ambiance proved ideal for a relaxed night out with friends or family. Don’t hesitate to give Sushi Station a try, you won’t regret it! 0.00
@@
SICO-Gen: I recently tried out Sushi Station’s revolving sushi bar and I must say, it was an amazing experience. The affordable pricing along with the wonderful krab mix and creamy sauce they offer is what really made it stand out from the rest. The menu was constantly updated with fresh and delicious options making it hard to choose just one. If you’re looking for a great sushi experience, Sushi Station is definitely worth a visit. 0.04
Table 6 presents the fake reviews generated by SICO-Gen and SICO-Para. The generated text shows high readability and fulfill the task’s goal, successfully mentioning all keywords and generating positive reviews of the specified object. The AI probability, denoted as $\mathcal{P}_{\text{AI}}$ in the table, is determined by the GPT3-D. More examples are shown in Appendix H.
6 Defensive Methods for SICO
6.1 Training Detectors with SICO Data
To defend against SICO, a straightforward approach is to train a detector on SICO text. We proposed an ensemble method, which utilizes two detectors trained separately to identify original AI-generated text and SICO text. The final AI probability for an input is calculated using the highest one from the two detectors, based on the assumption that both detectors should assign a low probability to human-written text. We avoided training a single detector with SICO text augmentation because our experiments indicated that this approach reduces the model’s ability to identify the original AI-generated text.
To test the effectiveness of this method, we obtain the SICO detector by fine-tuning a RoBERTa model with 5k SICO examples and 5k human-written examples. The human-written examples are sampled from WebText training set (Radford et al., 2019). The SICO examples are generated by using three different SICO-Para prompts to rewrite AI-generated text from the GPT2 output dataset (Solaiman et al., 2019). The SICO prompts used for detector training is entirely different from the one used for attack. We adopted the “GPT2-D” as the original AI-generated text detector. The results in Table 7 show that the ensemble approach significantly improves detection capabilities against SICO text, while maintaining the ability to identify the original AI-generated text. We believe this approach can be easily adapted to defend against other types of attacks, such as DIPPER, by training and incorporating detectors with adversarial attack.
Table 7: AUC scores of ensemble detector against normal AI-generated text and SICO text. “AI” refers to the detector’s AUC score on the original AI-generated text.
Dataset Detector AI SICO-Para SICO-Gen
Writing GPT2-D 0.848 0.332 0.099
[…]
QA GPT2-D 0.906 0.576 0.489
[…]
Review GPT2-D 0.952 0.264 0.619
[…]
6.2 Discussion of AI Detection Arms Race
Considering the evolution of AI detection techniques, the attackers and defenders are actually engaged in an arms race, where both sides continuously develop their own optimal strategies of variability and sophistication to overcome the opponents. We expect the newly proposed attack method will remain effective until the defender collects enough adversarial examples and trains a new detector on them. Then the attacker might develop a new technique to evade the most recent detector. This arms race presupposes that both sides are willing to share information like attack technique or detector API. If either side stops sharing, the race will come to a halt. Practically speaking, the defenders hold a superior position in this race, as they can restrict the access to detector, thereby preventing the attacker from improving their methods. Moreover, the defender, typically a large company, possesses more resources, including financial and computing power. The arms race represents the ultimate question in this field, which is complex and significant. Hence, we only provided a surface-level discussion. A more thorough investigation and experimentation on it are left for future research.
@@ 7 Conclusion
In conclusion, we have proposed a novel in-context learning approach, SICO, designed to guide LLMs in generating text that can effectively evade detectors. Our extensive experiments on evasion demonstrate the superior performance of SICO, which significantly reduces the detection capabilities of existing AI text detectors across three tasks. A comprehensive human evaluation shows SICO text can achieve human-level readability and task completion rates.
@@ 8 Ethics Statement
The intention of this paper is not to offer a potential method for evading AI-generated text detection systems. Instead, our aim is to raise awareness within the broader community about the vulnerabilities of existing AI-generated text detection systems to such technology. As many LLMs are public available and free to use, many people can adjust their prompt and generate text that evades these detectors. Given the ease of evasion illustrated in this study, these detectors are not robust yet. We hope the research community can stress test their detectors against text generated by carefully crafted prompt, and create more robust detectors in the future.
Besides presenting a potent attack technique, we also offer defensive methods against it. We believe future research will develop more sophisticated methods to enhance the robustness of AI detectors. To support the research in this field, we make our codes and data publicly available.
Acknowledgments
This work was supported by the National Key Research and Development Program of China under Grant 2022YFA1004102, and the Guangdong Major Project of Basic and Applied Basic Research (Grant No. 2023B0303000010).
@@ A.1.3 Word Substitution
We employ WordNet synsets to derive synonyms for given words. During optimization of in-context examples, we only substitute content words, namely nouns, verbs, adjectives, and adverbs. Furthermore, we part-of-speech (POS) tag of the synonym to ensure it aligns with the original word. For POS tagging, we utilize the Stanford POS Tagger (Toutanova et al., 2003). Additionally, to maintain fluency in the modified text after substitution, we employed a pretrained mask language model to exclude synonyms with low likelihood. In experiment we use RoBERTa-base model (Liu et al., 2019).
@@
Task Task instruction $p_{\text{task}}$
Writing Based on the description, complete an academic paragraph in human style writings:
@@
QA Based on the description, answer questions in human style writings:
@@
Review Based on the description, write a human review about the given object and keywords, with a specified sentiment:
@@
Paraphrase Based on the description, rewrite this in human-style writings:
@@ A.2.1 DIPPER
We choose the best evasion performance parameter setting from the original paper (Krishna et al., 2023), which is 60 for lexical diversity and 60 for re-ordering. And we set sampling temperature to 0.75, following the original implementation.
@@ A.2.2 Human prompt
We carefully design a paraphrase prompt based on the detection idea of GPTzero (Tian, 2023) and prompt shared online (Uploader, 2023), which distinguish the AI-generated content from Human-written by perplexity and burstiness, stated by its creator [^4] https://theconversation.com/we-pitted-chatgpt-against-tools-for-detecting-ai-written-text-and-the-results-are-troubling-199774. Perplexity is the concept raised in NLP field, which measures how well a language model predicts a text sample. A text with a lower perplexity score indicates that the language model is better at calculating the next word that is likely to occur in a given sequence. On the other hand, burtiness basically measures the variation between sentences, including sentence length and structures. The lower the values for these two factors, the more likely it is that a text was produced by an AI. Table 11 shows the prompt we designed.
@@ Appendix B Extra Experiments
B.1 Semantic preserving
We measure semantic similarity using t5-based sentence encoder (Ni et al., 2022), which leads in semantic text similarity task of MTEB benchmark (Muennighoff et al., 2023). Table 12 reports a comparison of the cosine similarity of text before and after paraphrasing by different methods.
@@ B.1 Semantic preserving
B.2 Stability of SICO
SICO is able to consistently construct effective detection evasion prompts, irrespective of the diversity in the initial AI-human text pairs and randomized samples drawn from the LLMs. This demonstrates the effectiveness of SICO in various initial conditions and settings, highlighting its applicability to diverse scenarios. Figure 3 presents the detection evasion performance of the best prompt at each step, denoted as $\mathcal{U}(p^{*})$ in Equation 1, where a higher value means higher evasion performance. The prompt is evaluated by 50 validation examples. We run three distinct experiment for each task to draw the plot. SICO successfully optimizes the initial prompt (at step 0) and achieves a high level of evasion performance across all three tasks, with different in-context examples.
@@ B.2 Stability of SICO
B.3 SICO performs better against more capable detectors
@@
B.4 SICO with different LLMs
To examine the effectiveness of SICO employing different LLMs, we adopted an additional experiment which employs SICO-Para on three different LLMs: WizardLM-13B Xu et al. (2023), GPT-3.5-turbo-0301, and GPT-4-0613 OpenAI (2022). We employ Chat-D as proxy detector. The results in Table 13 indicate that SICO with different LLMs maintains a high degree of detection evasion performance.
Table 13: The AUC scores of SICO using different LLMs. “-” indicates the performance without applying any evasion method. Symbol “*” represents that SICO uses GPT3-D as the proxy detector for prompt construction.
Dataset LLM GPT3-D* GPT2-D DetectGPT Log-Rank
Writing 0.908 0.848 0.834 0.914
WizardLM 0.571 0.582 0.510 0.481
GPT-3.5 0.239 0.332 0.149 0.147
GPT-4 0.157 0.260 0.263 0.277
QA 0.981 0.906 0.876 0.956
WizardLM 0.721 0.467 0.561 0.404
GPT-3.5 0.402 0.566 0.178 0.183
GPT-4 0.414 0.369 0.344 0.287
Review 0.925 0.952 0.808 0.982
WizardLM 0.707 0.402 0.570 0.542
GPT-3.5 0.465 0.264 0.270 0.300
GPT-4 0.408 0.307 0.427 0.382
@@ Appendix C Detectors
GPT-3.5 detector is trained on Human ChatGPT Comparison Corpus (HC3) dataset (Guo et al., 2023), which including answers generated by ChatGPT and human. English-version of HC3 dataset contains five splits: reddit-eli5, OpenQA, Wikipedia, medicine and finance. The base model is RoBERTa-base and we use the model that only take answers as input.
@@ C.3 DetectGPT
We use z-score implementation of DetectGPT and set sample number to 100 and replacement ratio to 0.3. The inner language model is GPT2-medium and the mask language model is t5-large (Raffel et al., 2020).
@@ C.6 OpenAI Detector
OpenAI detector is officially provided by OpenAI after the release of ChatGPT. Although it only offers a web interface, we adopted the API implementation from (Krishna et al., 2023), which uses “model-detect-v2” in the OpenAI API. Through reverse engineering of the website, we determined that the web interfaces indeed call this model.
@@ Appendix D Human Evaluation Details
For each text, we show two questions for the human annotator. In terms of readability, we present the human annotator five options, with the scale of 1 to 5. The higher the value, the more readable of the presented texts. The question is identical for three tasks. The actual question and options are presented in Table 14. For task completion rate, we design three task-specific questions, as show in Table 15. Figure 5 shows the interface of our annotation platform. We estimated that the evaluation time of each text ranges from 60 to 90 seconds. For the imperceptibility experiment, we present the annotators with text and a question to determine if ChatGPT generates the text. All annotators in the experiment used ChatGPT before.
@@
Appendix E Datasets
@@
Rev-Gen Write a review about given object and key words, with specified sentiment:
@@
Appendix F Evade Watermarking Detection
SICO-Para can also be utilized to evade watermark detection, similarly to paraphrase approach. The watermarking algorithm we applied was introduced by Kirchenbauer et al. (2023), which only requires access to the LLM’s logits at each time step to add watermarks. This algorithm operates in three steps:
@@ Appendix F Evade Watermarking Detection
In our experiments, we used the text generated by a watermarked GPT-2, provided by Krishna et al. (2023). We employed the GPT-3.5 detector as proxy detector for training. The AUC and and the detection accuracy associated with various paraphrasing methods are presented in Table 19, where the threshold is set to 2.2 for accuracy measurement.
The results reveal that SICO-Para significantly outperforms other paraphrase techniques in evading watermark detection. Notably, both the AUC score and detection accuracy of SICO-Para are lower than that of other methods.
@@
As we mentioned in Section 2, our work is focused on the vulnerability of training-based and statistical detectors. Therefore, we designed SICO to generate human-like text aimed at fooling these two types of detectors, which differentiate between human and AI-generated text based on language characteristics. However, as discussed in the previous section, watermarking methods distinguish human and AI-generated text by adding unnoticeable mathematical character in AI-generated text. Theoretically, any paraphrasing method that can significantly rewrite the original text will remove the watermark, thus evading the detector. For example, even the simplest baseline model, “GPT-Para”, proved to be effective at evading the watermark detection, reducing accuracy from 99% to 18%. Based on the aforementioned reasons, we put the watermark results in the appendix, though SICO-para can also help evade watermark detectors.
Appendix G Evasion Performance
G.1 ROC Curves
@@ G.2 Detection Accuracy
Task Feature
Writing Human tend to provide more specific details and facts, often including dates, numbers, and names. They also tend to use more complex sentence structures, including semicolons and parenthetical phrases, to convey additional information. However, their writing remains clear and concise, and they avoid unnecessary repetition or wordiness.
QA Human’s writing style is characterized by the provision of specific and technical information. They delve into the science behind a topic or the details of a process. They also challenges or adds nuance to the information, providing alternative explanations or clarifying misconceptions.
Review Human writings are generally shorter and more to-the-point. They often use bullet points or numbered lists to convey their thoughts. They tend to be more critical and focused on their own personal experiences. They also use informal language and occasionally include personal anecdotes or opinions.
Appendix H Examples of Text Generated by SICO
The examples of text generated by SICO across three tasks are presented in Tabel 22-24.
@@ Appendix H Examples of Text Generated by SICO
Question: Why does changing screen resolution (more specifically, a custom resolution) will harm your computer? $\mathcal{P}_{\text{AI}}$
AI: Changing the screen resolution, including a custom resolution, will not harm your computer. It may result in display issues or reduced image quality if the resolution is not compatible with your monitor’s capabilities. However, this can be resolved by simply reverting to the previous resolution or finding a suitable one. 0.99
SICO-Para: Don’t worry, changing your screen resolution won’t harm your computer, even if you decide to go custom. The only trouble is that it may not work well with your monitor and cause image problems or lesser quality. Easy solution though - just go back to the old resolution or find a matching one. 0.00
SICO-Gen: Using custom screen resolutions places a greater burden on system processing units, potentially leading to increased strain and overheating. Moreover, setting a resolution to unsupported display dimensions can put undue strain on graphics hardware, possibly causing irreparable damage. Adjusting to custom resolutions might, therefore, jeopardize the computer, leading to hardware malfunctions, degraded display output quality, and issues with other essential system components. 0.00
Question: How do they calculate how much calories you’ve burned after an exercise? $\mathcal{P}_{\text{AI}}$
AI: The number of calories burned during exercise is calculated by various factors such as body weight, the intensity of the exercise, the duration of the activity, and the person’s metabolic rate. Some popular fitness trackers and machines use algorithms and formulas based on these factors to estimate the number of calories burned. However, these numbers are only estimates and can vary depending on individual factors such as muscle mass, fitness level, and metabolic rate. 0.99
SICO-Para: If you wanna know how many calories you burn during exercise, there are a few things to keep in mind. Your weight, the intensity of your workout, and how long you’re doing it for all come into play, as does your metabolic rate. Popular fitness trackers and machines use fancy calculations based on all these factors to give you an estimate of how many cals you burned. But, listen up, folks - these are only estimates! Your muscle mass, fitness, and metabolism are gonna differ from person to person, so results will vary. Clear? 0.05
SICO-Gen: The calculation of calorie consumption requires assessment of parameters such as exercise type, duration, and intensity, as well as the individual’s body measurements. Several formulas exist to compute the total number of calories burned, including the Harris-Benedict equation, which computes resting metabolic rate and establishes parameters for calorie expenditure. Modern heart rate monitors and wearable fitness trackers utilise the latest algorithms to compute calorie consumption by taking into account the heart rate, age, and weight of the user. 0.00
@@
Object: hotel in Las Vegas Keywords: Penthouse, staff, casino, restaurants, excellent experience. Sentiment: Positive $\mathcal{P}_{\text{AI}}$
AI: I had an amazing experience staying at a penthouse in Las Vegas! The staff was incredibly friendly and accommodating, always going above and beyond to make my stay comfortable and enjoyable. The casino was also a great addition, providing endless entertainment and excitement. And when it came to dining options, the variety of restaurants available was impressive. Overall, staying at this hotel was an excellent experience that I won’t forget anytime soon. I highly recommend it to anyone looking for a luxurious and unforgettable Las Vegas getaway. 0.75
SICO-Para: Had an amazing stay at a penthouse in Las Vegas. The staff were incredibly friendly and made my stay comfortable and enjoyable. Casino provided endless entertainment and variety of restaurants impressive. Overall, excellent experience and highly recommend for luxurious and unforgettable Las Vegas getaway. 0.03
SICO-Gen: I had the most incredible experience staying at a luxurious penthouse in a Las Vegas hotel. The staff was exceptional and provided top service. The casino and restaurants on the premises were outstanding, making my stay all the more enjoyable. Overall, I highly recommend this hotel for an excellent experience in sin city. 0.06
Object: Tia Rosa restaurant Keywords: Mexican food, spicy, bland taste, sweet syrup, tacos. Sentiment: Negative $\mathcal{P}_{\text{AI}}$
AI: I recently visited Tia Rosa restaurant hoping to enjoy some authentic Mexican food. Unfortunately, my experience was far from satisfactory. The food was overly spicy, to the point where it overwhelmed any other flavors that may have been present. On the other hand, the tacos were disappointingly bland and lacked any real taste. To add insult to injury, the sweet syrup that was served with my meal just didn’t pair well with the savory dishes. Overall, my visit to Tia Rosa left me feeling disappointed and unsatisfied. 0.84
SICO-Para: Recently went to Tia Rosa restaurant looking for authentic Mexican food. However, experience was unsatisfactory. Food was overly spicy, overpowering other flavours. Tacos lacked taste. Sweet syrup served with meal did not complement savoury dishes. Overall left feeling unsatisfied and disappointed. 0.02
SICO-Gen: My visit to Tia Rosa restaurant was a complete letdown. I was expecting some spicy and flavorful Mexican food, but unfortunately, the dishes were quite bland. Even the tacos, which are usually a safe bet, were a disappointment. The only thing worthy of praise is the sweet syrup, but this alone cannot make up for the overall taste. I won’t be returning to Tia Rosa anytime soon. 0.04
@@
Writing Start: The code itself was patterned so that most control codes were together, and all graphic codes were together, for ease of identification. The first two columns (32 positions) were reserved $\mathcal{P}_{\text{AI}}$
@@
AI: The code itself was patterned so that most control codes were together, and all graphic codes were together, for ease of identification. The first two columns (32 positions) were reserved for control codes, while the remaining 94 columns (94 positions) were designated for graphic codes. This pattern allowed for efficient processing and recognition of the different types of codes. Additionally, the code was designed to be easily modifiable, with the ability to add or remove codes as needed. These features made the code a practical and versatile tool for programming and data processing tasks. 0.99